# More Tokens, Lower Precision: Towards the Optimal Token-Precision Trade-off in KV Cache Compression

### v1

@@ Abstract
As large language models (LLMs) process increasing context windows, the memory usage of KV cache has become a critical bottleneck during inference. The mainstream KV compression methods, including KV pruning and KV quantization, primarily focus on either token or precision dimension and seldom explore the efficiency of their combination. In this paper, we comprehensively investigate the token-precision trade-off in KV cache compression. Experiments demonstrate that storing more tokens in the KV cache with lower precision, i.e., quantized pruning, can significantly enhance the long-context performance of LLMs. Furthermore, in-depth analysis regarding token-precision trade-off from a series of key aspects exhibit that, quantized pruning achieves substantial improvements in retrieval-related tasks and consistently performs well across varying input lengths. Moreover, quantized pruning demonstrates notable stability across different KV pruning methods, quantization strategies, and model scales. These findings provide valuable insights into the token-precision trade-off in KV cache compression. We plan to release our code in the near future.
@@ 1 Introduction
As large language models (LLMs) have been widely used in various scenarios, such as document summarization (Zou et al., 2024), code completion (Roziere et al., 2023) and agent framework (Shridhar et al., 2020), there is a growing demand for models with larger context windows to handle more extensive inputs. As a result, GPT-4 (Achiam et al., 2023) have been extended to support 200k input tokens, and Gemini 1.5 (Reid et al., 2024) to 10M tokens. However, these powerful long-context capabilities come at the expense of significantly increased memory storage for the cached key and value (KV) states  (Waddington et al., 2013). Take Llama3-8B (Dubey et al., 2024) as an example, storing the KV cache of a single message with 100k tokens requires a high memory overhead of 20GB. Furthermore, as the decoding process heavily relies on the GPU memory bandwidth, the extensive KV cache also leads to dramatically increased decoding time (Fu, 2024).
To efficiently serve LLMs, various approaches have been proposed to compress KV cache (Pope et al., 2023). The predominant methods involve compressing the KV cache along two primary dimensions: token or precision. For the token dimension, KV pruning (or KV eviction) methods eliminate unimportant tokens to maintain a fixed KV cache size (Xiao et al., 2023; Zhang et al., 2024b; Ren and Zhu, 2024; Li et al., 2024). For the precision dimension, KV quantization technique reduces memory usage by approximating KV cache with lower precisions, like 8-bit or even lower (Sheng et al., 2023; Liu et al., 2024c; Hooper et al., 2024; Yang et al., 2024b). However, these existing works focuses only one dimension, either token or precision, leaving the trade-off between these two orthogonal factors largely under-explored.
Figure 1: More tokens, lower precision in KV cache leads to better performance.
@@
In this paper, we comprehensively investigate the token-precision trade-off in KV cache compression. First, we examine the feasibility of combining KV pruning and quantization under fixed budget. We demonstrate that storing more tokens in the KV cache with lower precision, which we call quantized pruning, can significantly enhance the long-context performance of LLMs. For example, with the same KV cache budget, storing 4×4$\times$4 × tokens in 4-bit precision outperforms storing 1×1$\times$1 × tokens in 16-bit precision across various downstream long-context tasks with various input length, as shown in Figure 1. Moreover, in extremely low-resource scenarios, quantized pruning effectively preserves performance, whereas relying solely on KV pruning or quantization often leads to a significant performance collapse.
Furthermore, we conduct in-depth analysis regarding token-precision trade-off from series of key aspects, including the impact on various downstream task types and input lengths, model scaling effect, ablation on quantization strategies and fine-grained exploration of layer-wise quantized pruning. From the extensive experiments, we observe that quantized pruning achieves substantial improvements in retrieval-related tasks and consistently performs well across varying input lengths. Moreover, quantized pruning demonstrates strong feasibility across different KV pruning methods, quantization strategies, and model scales, showcasing notable stability. The analysis on token-precision trade-off presents a more substantial compression potential compared to compressing along a single dimension. We believe that our findings could offer valuable insights for developing more effective KV compression strategies in future research.
@@ 2 Related Work
KV pruning compresses the KV cache along the token dimension by selectively removing unimportant tokens to reduce memory usage. Mainstream methods typically identify important tokens based on attention scores, as seen in (Liu et al., 2024b; Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024). Other methods use alternative factors such as initial tokens (Xiao et al., 2023), variance (Ren and Zhu, 2024), special tokens (Ge et al., 2024) or the L2 norm (Devoto et al., 2024) to determine token importance. Recent studies delve deeper into optimizing the allocation of KV cache memory budgets. Some explore KV cache budget allocation strategies across layers (Cai et al., 2024; Yang et al., 2024a), while other studies explore head-level KV cache budget allocation (Feng et al., 2024; Tang et al., 2024; Fu et al., 2024; Xiao et al., 2024).
@@
KV quantization compress KV cache from the precision dimension by storing KV cache using a reduced number of bits. FlexGen (Sheng et al., 2023) utilizes group-wise 4-bit quantization for both key and value cache. KIVI (Liu et al., 2024c) applies per-channel quantization on key cache and per-token quantization on value cache. KVQuant (Hooper et al., 2024) and CQ (Zhang et al., 2024a) use RoPE-related quantization, while KVQuant also preverses outliers without quantization. Atom (Zhao et al., 2024) reorders the outlier channels for fine-grained group quantization with mixed-precision. GEAR (Kang et al., 2024) uses low-rank approximation for quantization. QAQ (Dong et al., 2024) and MiKV (Yang et al., 2024b), inspired by the KV pruning methods, store discarded tokens using lower bit precision while retaining important tokens in full precision.
@@
Compressing KV cache from other dimensions typically requires modifying the model architecture, which usually necessitates additional training for adaptation. For the layer dimension, LCKV (Wu and Tu, 2024), CLA (Brandon et al., 2024) and MLKV (Zuhri et al., 2024) reduce memory usage by sharing the KV cache across adjacent layers. ShortGPT (Men et al., 2024) and DynamicSlicing (Dumitru et al., 2024) achieve compression by eliminating redundant layers. YOCO (Sun et al., 2024) changes the model structure and shares a single global KV cache across layers. For the head dimension, MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023) share the KV cache within each head groups. DeepSeek-v2 (Liu et al., 2024a) employs dimension-reduction techniques to compress all heads into a single low-rank vector. These lines of work is orthogonal to ours, as they can also be combined together.
3 Preliminaries
@@
Table 1: Feasibility of quantized pruned tokens on LongBench and Needle-in-a-Haystack with Llama-3-8B-Instruct as backbone model. We use four KV pruning methods to retain 128, 512 and 2048 tokens, and report the results of further quantization.
Figure 2: The token-precision trade-off under varying memory budgets on LongBench and NIAH. We report the results of SnapKV-based and PyramidKV-based quantized pruning on Llama-3 and Mistral-v0.2.
@@
Quantized pruning uses KV pruning methods to obtain $\bm{X}_{K}^{e}$ and $\bm{X}_{V}^{e}$ first, and then quantizes the preserved KV states $\bm{X}_{K}^{e}$ and $\bm{X}_{V}^{e}$ to Q($\bm{X}_{K}^{e}$) and Q($\bm{X}_{V}^{e}$) using various KV quantization methods in the prefill phase. In the decoding phase, the dequantized results from Q($\bm{X}_{K}^{e}$) and Q($\bm{X}_{V}^{e}$) are used to generate new tokens.
@@ 4 Experimental Setup
We assess the performances of quantized pruning using LongBench (Bai et al., 2024) dataset and Needle-in-a-Haystack (Kamradt, 2023) test. We also employ RULER (Hsieh et al., 2024), a dataset with different input length and diverse types of needles across 4 task categories, to better access the impact of input length in Section 6.1. More details can be seen in Appendix A.
@@
We use state-of-the-art open-weight LLMs, including Llama-3-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023). For scaling experiments in Section 6.2, we also test the performance of Llama-3.2-1B and Llama-3.2-3B (Dubey et al., 2024).
@@
We aim to comprehensively investigate the token-precision trade-off in KV cache compression. We report the ratio of compressed KV cache and the full KV cache for memory budge. For KV pruning, we employ PyramidKV (Cai et al., 2024) and SnapKV (Li et al., 2024), recognized as leading methods in diverse scenarios. We also include H2O (Zhang et al., 2024b) and StreamingLLM (Xiao et al., 2023) to evaluate the feasibility of quantized pruning in Section 5. For KV quantization, we adopt KIVI (Liu et al., 2024c) as the default method due to its stability and broad compatibility. Moreover, in Section 6.3, we examine the effects of quantization strategies from FlexGen (Sheng et al., 2023) and KVQuant (Hooper et al., 2024) for a comprehensive comparison. We use HQQ (Badri and Shaji, 2023) framework to perform quantization on KV cache. More details can be seen in Appendix B.
@@ 5 Optimal Token-Precision Trade-Off
In this section, we aim to find the optimal token-precision trade-off in KV cache compression. We frist examine the feasibility of combining KV pruning and quantization(Q1). Subsequently, we explore the best optimal allocation strategy between precision and token under varying memory budgets(Q2).
@@
We first evaluate the feasibility of quantized pruned KV cache as a prerequisite for exploring the token-precision trade-off. We use Llama-3-8B-Instruct and evaluate various KV pruning methods on the LongBench and NIAH. We report the results of quantizing the remaining tokens to different precision levels after applying KV pruning.
From Table 1, we observe that it is feasible to quantize pruned KV cache for a lowe compression rate.
For most KV pruning methods we evaluate, further quantizing the preserved tokens to as low as 4-bit precision results in minimal performance degradation, but quantizing to 8-bit precision shows negligible impact. However, reducing precision to 2-bit leads to a drastic performance decline across most KV pruning methods.
This observation holds consistently across different KV pruning methods and varying numbers of preserved tokens.
@@
Observing that KV pruning and KV quantization can be effectively combined, we further investigate that, given a fixed memory budget, how to balance the trade-off between number of preserved tokens and precision to achieve optimal performance. To this end, we evaluate the performance of quantized pruning using two leading KV pruning methods, SnapKV and PyramidKV, across different memory budgets on LongBench and NIAH.
Specifically, we compare three configurations with approximately equivalent memory usage: 1) Using standalone KV pruning to retain 1×1$\times$1 × tokens in 16-bit precision. 2) Quantized pruning by retaining 2×2$\times$2 × tokens in 8-bit precision. 3) Quantized pruning by retaining 4×4$\times$4 × tokens in 4-bit precision.
@@ 6 Further Analysis
Table 2: The token-precision trade-off in different task types. We report the results of 6 task types in LongBench and 8k subset of RULER. We use PyramidKV-based quantized pruning.
6.1 Impact on Downstream Tasks
@@
To further investigate the token-precision trade-off in different task types, we evaluate PyramidKV-based quantized pruning on six task types from LongBench and the 8K subset of the RULER dataset. We use PyramidKV with 512 retained tokens as the baseline, and explore the token-precision trade-off under this fixed memory budget, as this setting exhibits minimal performance differences across three precision levels, making it easier to assess the impact of task types.
As illustrated in Table 2, we observe that the performance of quantized pruning is remarkably consistent across different task types. Specifically, lower precision, which retains more tokens in KV cache, leads to substantial performance improvements in the RULER task, which heavily relies on retrieving content from the input. Tasks with high retrieval demands, such as Summarization and Single-Doc QA, also show noticeable gains with quantized pruning, particularly when 4×4$\times$4 × tokens are preserved at 4-bit precision.
For tasks requiring more reasoning rather than intensive retrieval, such as Code Completion, Synthetic and Multi-Doc QA, the benefits of trading precision for more tokens are less pronounced. In these cases, storing fewer tokens with higher precision generally performs better. For example, using 1024 tokens in 8-bit precision achieves the hightest score of 58 in Code task with Llama-3.
@@
To evaluate the token-precision trade-off across various input lengths, we conduct experiments on subsets with different input length of the RULER dataset. Additionally, we analyze LongBench by grouping its data based on input length. The results is shown in Figure 3 and more detailed information can be found in Appendix A.
Our observations are as follows: quantized pruning consistently outperforms standalone KV eviction methods across various input length, regardless of the models and task types. Within the same dataset, scores decrease as input length increases; however, the relative differences among different compression methods remain similar across varying input lengths. Moreover, quantized pruning achieves significant performance improvements across all input lengths for retrieval demanded tasks like RULER.

### v2

@@ Abstract
As large language models (LLMs) process increasing context windows, the memory usage of KV cache has become a critical bottleneck during inference. The mainstream KV compression methods, including KV pruning and KV quantization, primarily focus on either token or precision dimension separately. However, these works leaving the trade-off between these two orthogonal dimensions largely under-explored. In this paper, we comprehensively investigate the token-precision trade-off in KV cache compression. Experiments demonstrate that storing more tokens in the KV cache with lower precision, a strategy we term quantized pruning, can significantly enhance the long-context performance of LLMs. In-depth analysis of the token-precision trade-off across key aspects demonstrates that, quantized pruning achieves substantial improvements in retrieval-related tasks and consistently performs well across varying input lengths. Furthermore, quantized pruning demonstrates notable stability and effectiveness across different KV pruning methods, quantization strategies, and model scales. These findings offer valuable insights into optimizing KV cache compression through balanced token-precision trade-off strategies. Our code is available at https://github.com/zhzihao/QPruningKV.
@@ 1 Introduction
Current long-context Large Language Models (LLMs) heavily depend on the Key-Value (KV) cache mechanism, which stores intermediate keys and values during text generation to avoid redundant computations (Waddington et al., 2013). As the input sequence length increases, the KV cache size expands proportionally, leading to substantial memory overhead (Achiam et al., 2023; Reid et al., 2024). Take Llama3-8B (Dubey et al., 2024) as an example, storing the KV cache of a single sequence with 100k tokens necessitates a substantial memory overhead of 20GB — exceeding the model’s parameter memory of approximately 14GB. Moreover, since the decoding process is highly dependent on GPU memory bandwidth, the large KV cache also results in a significant increase in decoding time (Fu, 2024). As a result, efficiently compressing the KV cache has become a critical challenge in advancing LLM development (Pope et al., 2023).
To optimize KV cache memory usage, two primary approaches, focusing on token and precision dimensions, are widely adopted: KV pruning and KV quantization. For the token dimension, KV pruning methods reduce the KV cache footprint by discarding less salient tokens, thereby maintaining a fixed cache size (Xiao et al., 2023; Zhang et al., 2024b; Ren and Zhu, 2024; Li et al., 2024). For the precision dimension, KV quantization methods reduce memory usage by approximating KV cache with lower precisions, like 4-bit or even lower (Sheng et al., 2023; Liu et al., 2024c; Hooper et al., 2024; Yang et al., 2024b). Given the orthogonal nature of token and precision dimensions in KV cache, it is natural to consider integrating KV pruning and KV quantization to balance the number of tokens and their precision for better performance. For a fixed cache budget, lower precision in KV quantization would allow for the retention of more tokens in KV pruning. This raises the important questions: Is such an integration effective? How can we achieve an optimal trade-off between pruning and quantization? To the best of our knowledge, prior research has not explored this interplay.
@@
To address the aforementioned questions, in this paper, we comprehensively explore the token-precision trade-off in KV cache compression. Through meticulous experimentation, we for the first time present the feasibility of integrating KV pruning and KV quantization. Specifically, we uncovered an intriguing finding: storing more tokens in the KV cache with lower precision, a strategy we term quantized pruning, can significantly enhance the long-context performance of LLMs. For instance, storing 4×4$\times$4 × tokens in 4-bit precision outperforms storing 1×1$\times$1 × tokens in 16-bit precision across various KV cache budget. Notably, in extremely low-resource scenarios, quantized pruning effectively preserves performance, whereas relying solely on KV pruning or quantization often leads to a significant performance collapse. Furthermore, we conduct in-depth analysis of the token-precision trade-off across critical aspects, including task type, input length, model scaling, quantization strategies, and layer-wise configurations. Extensive experiments reveal that quantized pruning achieves considerable gains in retrieval tasks and maintains robust performance across varying input lengths. Moreover, quantized pruning demonstrates strong stability across different KV pruning methods, quantization strategies, and model scales.
In summary, our contributions are as follows:
  • We are the first to comprehensively explore the integration of KV pruning and quantization and propose the novel strategy of quantized pruning to enhance the long-context capabilities of LLMs.
  • An important finding is revealed: storing more tokens with lower precision significantly outperforms standalone KV pruning or KV quantization methods under various KV cache budgets, highlighting the importance of balancing token count and precision in KV cache compression.
  • Extensive experiments have been conducted on exploring token-precision trade-off across various critical aspects, including task types, input lengths, model scales, and quantization strategies, providing valuable insights for optimizing KV cache compression.
@@ 2 Related Work
KV pruning compresses the KV cache along the token dimension by selectively discarding less salient tokens to reduce memory usage. Mainstream methods typically identify salient tokens based on attention scores, as seen in (Liu et al., 2024b; Zhang et al., 2024b; Oren et al., 2024; Li et al., 2024). Other methods use alternative factors such as initial tokens (Xiao et al., 2023), variance (Ren and Zhu, 2024), special tokens (Ge et al., 2024) or the L2 norm (Devoto et al., 2024) to determine token importance. Recent studies delve deeper into optimizing the allocation of KV cache memory budgets. Some explore KV cache budget allocation strategies across layers (Cai et al., 2024; Yang et al., 2024a), while other studies explore head-level KV cache budget allocation (Feng et al., 2024; Tang et al., 2024; Fu et al., 2024; Xiao et al., 2024).
@@
KV quantization compress KV cache from the precision dimension by storing KV cache using a reduced number of bits. FlexGen (Sheng et al., 2023) utilizes group-wise 4-bit quantization for both key and value cache. KIVI (Liu et al., 2024c) applies per-channel quantization on key cache and per-token quantization on value cache. KVQuant (Hooper et al., 2024) and CQ (Zhang et al., 2024a) use RoPE-related quantization, while KVQuant also preverses outliers without quantization. Atom (Zhao et al., 2024) and SKVQ (Duanmu et al., 2024) reorders the outlier channels for fine-grained group quantization with mixed-precision. QAQ (Dong et al., 2024) and MiKV (Yang et al., 2024b), inspired by KV pruning methods, store discarded tokens from KV pruning methods using lower bit precision while retaining salient tokens in full precision. However, their memory usage scales proportionally with context length. In contrast, quantized pruning, fundamentally based on KV pruning, theoretically enables compression to a pre-defined memory budget.
@@
Compressing KV cache from other dimensions typically requires modifying the model architecture, which usually necessitates additional training for adaptation. For the layer dimension, LCKV (Wu and Tu, 2024), CLA (Brandon et al., 2024) and MLKV (Zuhri et al., 2024) reduce memory usage by sharing the KV cache across adjacent layers. ShortGPT (Men et al., 2024) and DynamicSlicing (Dumitru et al., 2024) achieve compression by eliminating redundant layers. YOCO (Sun et al., 2024) changes the model structure and shares a single global KV cache across layers. For the head dimension, MQA (Shazeer, 2019) and GQA (Ainslie et al., 2023) share the KV cache within each head groups. DeepSeek-v2 (Liu et al., 2024a) employs dimension-reduction techniques to compress all heads into a single low-rank vector. These lines of work are orthogonal to ours, and theoretically, they could be combined with our method. However, it is important to note that these methods typically necessitate adjustments at the model architecture or require additional fine-tuning, while quantized pruning can be directly employed without any fine-tuning.
3 Methods
@@
Ada-KV 87.7 88.2 81.0 11.1 98.6 98.6 97.4 11.9 100 100 100 27.3
HeadKV 98.6 98.5 98.4 9.1 100 99.9 99.7 11.5 100 100 100 27.6
Table 1: Feasibility of quantized pruned tokens on LongBench and Needle-in-a-Haystack with Llama-3-8B-Instruct as backbone model. We use six different KV pruning methods to retain 128, 512 and 2048 tokens, and report the results of further quantization. KIVI is used as the default quantization method, except for Ada-KV and HeadKV, which adopt FlexGen. As they retain different tokens per head, making them difficult to be compatible with KIVI. The best results for each token-precision setting are in bold, the second best results are underlined.
Figure 1: The token-precision trade-off under varying memory budgets on LongBench and NIAH. We report the results of SnapKV-based and PyramidKV-based quantized pruning on Llama-3 and Mistral-v0.2. We compare three configurations with approximately equivalent memory usage: 1) Using standalone KV pruning to retain 1×1$\times$1 × tokens in 16-bit precision. 2) Quantized pruning by retaining 2×2$\times$2 × tokens in 8-bit precision. 3) Quantized pruning by retaining 4×4$\times$4 × tokens in 4-bit precision. Quantized pruning, which preserves more tokens at a lower precision, consistently outperforms standalone KV pruning methods across various budgets.
@@
Quantized pruning uses KV pruning methods to obtain $\bm{X}_{K}^{e}$ and $\bm{X}_{V}^{e}$ first, and then quantizes the preserved KV states $\bm{X}_{K}^{e}$ and $\bm{X}_{V}^{e}$ to Q($\bm{X}_{K}^{e}$) and Q($\bm{X}_{V}^{e}$) using KV quantization methods in the prefill phase. In the decoding phase, the dequantized results from Q($\bm{X}_{K}^{e}$) and Q($\bm{X}_{V}^{e}$) are used to generate new tokens.
@@ 4 Experimental Setup
We evaluate the performance of quantized pruning using the LongBench benchmark (Bai et al., 2024) and Needle-in-a-Haystack test (Kamradt, 2023). To better distinguish the performance impacts related to input lengths and layer-wise configurations (Sections 6.1 and 6.4), we further utilize RULER (Hsieh et al., 2024), a dataset with different input length and diverse types of needles across 4 task categories. Detailed information on these datasets can be found in Appendix A.
@@
In our primary evaluations, We employ state-of-the-art open-weight LLMs, specifically Llama-3-8B-Instruct (Dubey et al., 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023). For scaling experiments in Section 6.2, we additionally assess the performance of Llama-3.2-1B, Llama-3.2-3B (Dubey et al., 2024) and Llama-3-70B (Dubey et al., 2024).
@@
Our experiments are designed to comprehensively investigate the token-precision trade-off in KV cache compression. We quantify memory budget by reporting the compression ratio of the KV cache relative to the full, uncompressed KV cache. For KV pruning, we utilize PyramidKV (Cai et al., 2024) and SnapKV (Li et al., 2024) as representative state-of-the-art methods. To assess the feasibility of quantized pruning (Section 5), we also include StreamingLLM (Xiao et al., 2023), H2O (Zhang et al., 2024b), Ada-KV (Feng et al., 2024), and HeadKV (Fu et al., 2024). For KV quantization, we adopt KIVI (Liu et al., 2024c) as our primary method due to its stability and broad compatibility. Furthermore, in Section 6.3, we evaluate quantization strategies from FlexGen (Sheng et al., 2023) and KVQuant (Hooper et al., 2024) for a comprehensive comparative analysis. Additional setup details are provided in Appendix B.
@@ 5 Optimal Token-Precision Trade-Off
In this section, we aim to find the optimal token-precision trade-off in KV cache compression. We first examine the feasibility of integrating KV pruning and quantization(Q1). Subsequently, we explore the best optimal allocation strategy between precision and token under varying memory budgets(Q2).
@@
We first evaluate the feasibility of integrating KV pruning and quantization as a prerequisite for exploring the token-precision trade-off. We employ Llama-3-8B-Instruct and evaluate a range of KV pruning methods on the LongBench and NIAH. We report the results of quantizing the remaining tokens to different precision levels after applying KV pruning.
From Table 1, we observe that it is feasible to quantize pruned KV cache for a lower compression rate. For most KV pruning methods we evaluate, further quantizing the preserved tokens to as low as 4-bit precision results in minimal performance degradation, while quantizing to 8-bit precision shows negligible impact. However, reducing precision to 2-bit leads to a drastic performance decline across most KV pruning methods. This observation holds consistently across different KV pruning methods and varying numbers of preserved tokens.
Notably, head-level KV allocation methods such as Ada-KV and HeadKV, which are incompatible with KIVI, perform well under high precision and a low number of preserved tokens. However, when precision is reduced to as low as 2-bit, these methods experience a more pronounced performance degradation compared to KV pruning methods compatible with KIVI, such as SnapKV and PyramidKV. This observation underscores the critical importance of considering compatibility with KV quantization techniques in the development of future KV pruning strategies.
@@
To conclude, KV pruning can be effectively integrated with KV quantization at a precision level of 4-bit without substantial performance degradation.
@@
Observing that KV pruning and KV quantization can be effectively integrated, we further investigate that, given a fixed memory budget, how to balance the trade-off between number of preserved tokens and precision to achieve optimal performance. To this end, we evaluate the performance of quantized pruning using two leading KV pruning methods, SnapKV and PyramidKV, across different memory budgets on LongBench and NIAH.
@@ 6 Further Analysis
Table 2: The token-precision trade-off in different task types. We report the results of 6 task types in LongBench and 8k subset of RULER. We use PyramidKV-based quantized pruning. The best results for each task type are in bold.
6.1 Impact on Task Types and Input Lengths
@@
To further investigate the token-precision trade-off in different task types, we evaluate PyramidKV-based quantized pruning on six task types from LongBench and the 8K subset of the RULER dataset. We use 1/16 KV cache budget, and explore the token-precision trade-off under this fixed memory budget, as this setting exhibits minimal performance differences across three settings, facilitating a more direct comparison of performance across task types.
As illustrated in Table 2, we observe that the performance of quantized pruning is remarkably consistent across different task types. Specifically, lower precision, which retains more tokens in KV cache, leads to substantial performance improvements in the RULER task, which heavily relies on retrieving contents from the input. Tasks with high retrieval demands, such as Summarization and Single-Doc QA, also show noticeable gains with quantized pruning, particularly when 4×4$\times$4 × tokens are preserved at 4-bit precision.
For tasks requiring more reasoning rather than intensive retrieval, such as Code Completion, Synthetic and Multi-Doc QA, the benefits of trading precision for more tokens are less pronounced. In these cases, storing fewer tokens with higher precision generally performs better. For example, using 1024 tokens in 8-bit precision achieves the hightest score of 58 in Code task.
@@
To evaluate the token-precision trade-off across various input lengths, we conduct experiments on subsets with different input length of the RULER dataset. Additionally, we analyze LongBench by grouping its data based on input length, more detailed information can be found in Appendix A. The results are shown in Figure 2.
Our observations are as follows: quantized pruning consistently outperforms standalone KV eviction methods across various input lengths, regardless of the models and task types. Within the same dataset, scores decrease as input length increases; however, the relative differences among different compression methods remain similar across varying input lengths. Moreover, quantized pruning achieves significant performance improvements across all input lengths for retrieval demanded tasks like RULER.